# Large Language Model Aided QoS Prediction for Service Recommendation

### v1

@@ V Conclusion
In this paper, we propose large language model aided QoS prediction (llmQoS) model. It utilizes LLMs to extract features from descriptive sentences constructed based on attributes of web users and services. Then the LLM features are combined with ID based embedding features for QoS prediction through a multi-layer perceptron network. Despite its concise network architecture, llmQoS demonstrates the ability of accurate QoS prediction under data sparsity condition, and outperforms comparable baseline models consistently on the WSDream dataset. We also show that the benefits of adding LLM features is general, and larger LLMs achieves better performance. In the era of explosive development of LLMs, we hope our paper can inspire more researchers in the field of service recommendation to conduct research in this direction.

### v2

@@ V Conclusion
In this paper, we have proposed the large language model aided QoS prediction (llmQoS) model. It utilizes LLMs to extract features from descriptive sentences constructed based on attributes of web users and services. Then the LLM features are combined with ID based embedding features for QoS prediction through a multi-layer perceptron network. Despite its concise network architecture, llmQoS achieves accurate QoS prediction under data sparsity condition, and outperforms comparable baseline models consistently on the WSDream dataset. We have also shown that the benefits of adding LLM features is general, and larger LLMs achieve better performance. In the era of explosive development of LLMs, we hope our paper can inspire more researchers in the field of service recommendation to conduct research in this direction.